OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = negative 2 over 3x + 3.

OpenStudy (danjs):

recall what parallel lines have in common

OpenStudy (danjs):

The Line passes through point (7,3) and has the same slope as y = -(2/3)x + 3

OpenStudy (danjs):

point slope form of a line y - y1 = m*(x-x1) where given point (x1,y1) = (7,3) and slope m= -2/3

OpenStudy (anonymous):

ok so I got y-3=-2/3 * (x-7) grr I used to know how to solve these equations freshman year but it all faded :( @DanJS

OpenStudy (anonymous):

@DanJS

Nnesha (nnesha):

yesh that's right

Nnesha (nnesha):

now solve for y

Nnesha (nnesha):

he is busy i guess just want to tell you that is right keep working onthat

OpenStudy (anonymous):

how do do that?

OpenStudy (anonymous):

I*

Nnesha (nnesha):

 distribute bracket by -2/3